# Automating Governing Knowledge Commons and Contextual Integrity (GKC-CI) Privacy Policy Annotations with Large Language Models

### v2

@@ 5.4. Qualitative Error Analysis
In order to better understand the errors made by GPT 3.5TPE_25ep, we performed qualitative coding on the 37 match errors for positive examples produced by the model, i.e., match errors where neither the ground truth label nor the model completion was “N/A”. This served two purposes.
First, we identified cases where labels were mistakenly labeled as errors, specifically where the model annotation was semantically equivalent, albeit syntactically different from the ground truth.
Second, we conducted a detailed analysis of the model’s performance, examining the factors contributing to both its strengths and weaknesses. This provided more confidence in overall model performance.
Identifying trends in these errors offers valuable insights into the model’s behavior. Although these insights may not explain why an error occurs (LLM models have notoriously poor explainability (Druce et al., 2021; Barredo Arrieta et al., 2020)), they help users become aware of the model’s limitations. This enables users to use the model more strategically, leveraging its strengths and compensating for its weaknesses, thereby enhancing its overall effectiveness.
@@ 5.4.1. Qualitative Coding
To ensure the reliability and consistency of the coding process, two expert coders initially met to collaboratively develop a comprehensive codebook consisting of ten codes: three parent codes and seven child codes. The full codebook can be seen in Appendix G.
After joint codebook creation, each coder independently coded all match errors produced by GPT 3.5TPE_25ep. Once the coding was complete, we computed inter-coder reliability and found a high level of agreement between the two coders with a Cohen’s kappa score of .87 (Cohen, 1960). The results of the qualitative coding are visualized in Figure 5 and detailed further below.
@@ 5.4.2. Semantic Equivalence
The code “Semantic Equivalence”, which is the sole code of the Semantic Equivalence category, was the most prevalent in our error analysis—accounting for 19/37 (51%) errors. The following two examples demonstrate what this error looks like. The text in quotation is the expert annotation, while the underlined sections are the model’s response:
@@
Figure 5. Breakdown by category of the various types of errors found from our qualitative analysis.
5.4.3. Incorrect Expert Labels
The category Expert Labels Is Wrong had nine examples (24%). Occasional expert mis-annotations are expected for a task of this complexity. We are encouraged that there were relatively few examples in this category, supporting the quality of our ground truth. Importantly, for the examples in the “Expansive Ground Truth” (4/9, 44%) and “Partial Ground Truth” (3/9, 33%) child codes, the model performed the task more correctly than the expert annotator—either by omitting superfluous words included in the expert annotation or including necessary ones the expert annotator missed.
Consider the following example “expansive ground truth” annotation. The text in quotation is the expert annotation, while the crossed sections are what the model correctly excluded.
@@ 5.4.3. Incorrect Expert Labels
Conversely, the following example “partial ground truth” annotations show the expert annotation in quotations, while the underlined words are what the model correctly choose to additionally include in its response.
@@
Combined, “Expansive Ground Truth” and “Partial Ground Truth” codes represent only 9 of the 37 errors (24%); however, we emphasize this finding as particularly exciting because they demonstrate that the model can identify precise annotations for the requested parameter. In other words, the model’s ability has surpassed that of our highly-trained human annotators for these situations.
@@ 5.4.4. True Model Errors
The category “Completion Errors” accounted for 9 out of 37 coded examples (24%). Notably, a significant majority of these errors (6 out of 9, 67%) fall under the “meaningful subset” child code. A “meaningful subset“ annotation included a segment of the correct response, but missed words that altered the meaning of the annotation. In the following example, the text in quotation is the expert annotation while the underlined sections are the model’s response:
@@
Observe that the model’s completion is a sentence fragment, regardless of the parameter specified. Many of these fragments do encompass enough of the correct answer for someone well-versed in CI-GKC to grasp the intended annotation. However, we consider these incomplete responses to be incorrect even though the model’s answer is a meaningful subset of the correct response.
Furthermore, the code “Completion is wrong” contains the remaining elements Completion Is Wrong category, comprising the final 3 out of 9 codes (33%). All of these responses labeled text from within the sentence with no relation to the actual GKC-CI parameter. For example the model incorrectly labels the underlined part of the following sentence as an attribute:
@@
Our qualitative analysis offers a comprehensive view of the model’s performance, detailing both its strengths and limitations. Appreciably more than half of the purported match errors can be attributed to “Semantic Equivalence”, “Expansive Ground Truth”, and “Partial Ground Truth.” These combined child codes constitute 27 examples, resulting in a $0.29$ percent increase in the number of correct annotations overall. This implies that our benchmarking metrics for model accuracy serves as a conservative estimate of model performance. In essence, while our model displays commendable performance, there is still room for improvement to address model errors.
@@ 5.5. Training Data Review
Motivated by the nine expert labeling errors we uncovered in Section 5.4.3, we had two different CI experts review a random sample of 5% of our training data. These experts agreed with the training data on 90.57% of the sample, with the most common disagreement being the identification of occasional attribute parameters that were missed in the original annotations. The original annotators spent several months annotating thousands of parameters, so it is unsurprising that a few were missed. Furthermore, governing knowledge commons and contextual integrity are the topic of active research and subject to different interpretive nuances by different members of the community, which is reflected in different annotations. This training data review validates our results, as our best performing model’s 90.65% accuracy is close to the inter-expert annotation agreement, suggesting that the model approaches human performance on the task.
@@ 6. Limitations
Second, the LLMs we use do not have information about the input text outside of the context window. This is unlike a human lawyer, who could review the entire document while identifying information flow descriptions. Fortunately, the information needed to identify GKC-CI parameters typically appears in the text immediately surrounding the parameter itself. The decreased accuracy we observed when using a longer context window (Section 5.3.2, Appendix C) supports our use of one sentence contexts.
Third, we do not perform coreference resolution, i.e., identifying when different sections of text refer to the same real world entity. For example, a sentence may refer to the website user multiple times, with each instance receiving a GKC-CI parameter annotation (e.g., “sender”). Although our model would accurately annotate each mention of the sender, it would not automatically recognize that these annotations refer to the same legal entity. If that were necessary for a downstream application, another model would be required. Since coreference resolution is an active research problem in its own right, and is independent from GKC-CI parameter labeling, we leave this for future development.
@@ 7. Example Applications
We applied our GPT 3.5TPE_25ep model to 456 privacy policies from the Princeton-Leuven Longitudinal Corpus of Privacy Policies (Amos et al., 2021) to demonstrate the type of analyses enabled by GKC-CI annotation at scale. This dataset contains over 1 million privacy policies from over 100,000 companies spanning more than two decades, making it an ideal data source. However, we note that the primary contribution of our project remains the LLM training and evaluation (Sections 45). This section is not meant to provide a comprehensive analysis of policies in the Princeton-Leuven dataset. Rather, we intend the following examples to inspire future work using our LLM annotation method and the annotated policies we provide. All 456 policies we annotated for the following analyses are publicly available at the GitHub repository for this paper. [^8] https://github.com/JakeC007/Automated_GKC-CI_Privacy_Policy_Annotations
@@ 7.1. Longitudinal Privacy Policy Analysis
First, we chose 8 prominent companies and organizations [^9] Facebook, The New York Times, GitHub, Buzzfeed, Google, Bank of America, Electronic Frontier Foundation (EFF), and the National Science Foundation (NSF) representing a variety of sectors, including “big tech,” news, entertainment, finance, and government. This variety is useful because each sector has a unique approach to data collection, user engagement, and compliance with privacy regulations. Furthermore, these websites have undergone varying levels of public scrutiny. For instance, while Facebook and Google have faced major privacy debates, leading to numerous changes in their privacy policies, entities like nsf.gov operate under distinct governmental standards. This list also highlights geographic diversity concerning headquarters and user base, with some companies primarily serving U.S. audiences, like Bank of America, while others have a global reach, necessitating compliance with various international privacy laws per the Brussels effect (Bradford, 2012).
For each of these companies and organizations, we used our model to annotate one privacy policy from every year that the company or organization appears in the dataset (128 policies total). The number of parameters in the policies of each company or organization over time are presented in Appendix H (Figure 17 and Tables 56).
@@ 7.2. Cross-Industry Privacy Policy Analysis
Figure 6 shows the fifteen policies with the highest variance of parameter type percentages. This includes a policy from apache.org with relatively few aim parameters and a policy from mozilla.org with relatively few attribute parameters. A quick review of the mozilla.org policy shows why – it defines all data covered by the policy in one sentence: “For us, ‘personal information’ means information which identifies you, like your name or email address.” The policy does not say which particular forms of personal information are connected to the descriptions of information use in the rest of the policy, leaving it ambiguous whether e.g., what specific customer data might be shared with other entities for “processing or providing products and services.” While this type of high-level analysis doesn’t necessarily imply the existence of policy ambiguities, it suggests that policies with high parameter type variance are promising candidates for a detailed evaluation through the lens of the GKC-CI framework. The parameter percentages for all 164 privacy policies are provided in Tables 78 in Appendix I.
@@
We next calculated the ratio of annotated GKC-CI parameters to the number of sentences in each policy. This provides a metric of the “density” of information transfer descriptions in the policy. Figure 7 shows these data for the 15 privacy policies with the highest ratio of annotated parameters to sentences. The ratios for all 164 privacy policies are provided in Tables 910 in Appendix J. The top 15 policies include those from content distribution networks (b-cdn.net), web component frameworks (ampproject.org), Facebook, Microsoft (bing.com), Apple (icloud.com), Zoom, and social media websites (linkedin.com, snapchat.com, t.co, tumblr.com), among others.
@@ 7.3. GKC-CI Annotation Visualizer
We have developed a visualizer tool designed to enhance the interpretability of our LLM’s privacy policy annotations. This tool was created to empower users to better understand the complex data flows and governing principles that underpin privacy policies by visually distinguishing between different parameter tags. By highlighting various aspects of the policy in different colors, the visualizer makes the described information flows more accessible, allowing users to quickly grasp how their data may be used and protected under the terms of the policy.
The tool functions as a local Python script that processes both the raw text of the privacy policy and the output log from the LLM. It begins by matching specific text segments from the policy with those used in the LLM prompts. The visualizer then highlights these segments in different colors based on the associated annotations, making it easy for users to see which parts of the text correspond to specific GKC-CI parameters. This highlighted text is presented in a user-friendly GUI window, where users can scroll through and interact with the augmented policy. Additionally, it generates an output file that explains the highlighted segments, helping users understand the implications of the annotations.
To illustrate the capabilities of the visualizer, we applied it to two versions of Facebook’s privacy policy from 2015 and 2019. Comparing visualized excerpts of these versions (Appendix K) effectively highlights how data practices and governing principles have evolved. For instance, there are more GKC-CI parameters in the 2019 excerpt than the 2015 excerpt, particularly more aims and consequences. The new policy does a better job informing users about the repercussions of the described data collection. The visualizer’s color-coded presentation makes it easier to understand how the policy has changed and will facilitate future research using GKC-CI annotation.
@@ 9. Conclusion
This paper demonstrates that high-accuracy annotation of contextual integrity (CI) and governing knowledge commons (GKC) parameters in privacy policies can be achieved using LLMs. We ultimately find that GPT 3.5TPE_25ep had the best performance, with an accuracy of 90.65% for exact string matches. While we find that the proprietary LLMs outperformed open-source models, we report some valuable findings for researchers interested in performing LLM application studies. Namely, 1) that LLM size must be considered in context to model family; smaller, aligned models be the most economical choice, and 2) that library defaults are likely to introduce confounds and should be checked.
We demonstrate the usefulness of our fine-tuned model by annotating 456 privacy policies from the Princeton-Leuven Longitudinal Corpus of Privacy Policies (Amos et al., 2021). We demonstrate that large-scale GKC-CI annotation can be an effective tool for data exploration, highlighting changes in parameter frequency over time, policies with relatively high variances across parameter type percentages, and policies with relatively high ratios of parameters to sentences. This facilitates automated review of privacy policy updates to identify meaningful changes in information flow descriptions, potentially with normative implications. We have made our model training code, training and testing data, annotation visualizer, and all annotated policies publicly available [^10] https://github.com/JakeC007/Automated_GKC-CI_Privacy_Policy_Annotations in the hope that this work motivates future use of GKC-CI parameter annotation.
Acknowledgements
@@ Appendix A Ground Truth Details
Table 3. Number of labeled parameters in ground-truth GKC-CI annotations of 16 privacy policies from popular websites and e-learning services.
@@ Appendix B Brat Annotation Legend
Figure 9. Legend in the customized Brat Annotation tool for expert annotators to use as reference.
Appendix C Performance On Paragraph Delineated Data As Epochs Vary
Figure 10. The dotted orange line (“FT-Paragraph Accuracy”) is the test-set accuracy of GPT 3.5TPE trained on paragraph delineated data at 1, 5, 10, 25, and 50 epochs. The solid blue line (“FT-Sentence Accuracy”) is the test-set accuracy of GPT 3.5TPE trained on sentence delineated data at 1, 5, 10, 25, and 50 epochs. An accurate annotation is one that is an exact string match.

### v3

@@ 5.4. Qualitative Error Analysis
In order to better understand the errors made by GPT 3.5TPE_25ep, we performed qualitative coding on the 37 match errors for positive examples produced by the model, i.e., match errors where neither the ground truth text nor the model completion was “N/A”. This served two purposes.
First, we identified cases where model outputs were mistakenly annotated as errors, specifically where the model annotation was semantically equivalent, albeit syntactically different from the ground truth. Second, we conducted a detailed analysis of the model’s performance, examining the factors contributing to both its strengths and weaknesses. This provided more confidence in overall model performance.
Identifying trends in these errors offers valuable insights into the model’s behavior. Although these insights may not explain why an error occurs (LLM models have notoriously poor explainability (Druce et al., 2021; Barredo Arrieta et al., 2020)), they help users become aware of the model’s limitations.
@@ 5.4.1. Qualitative Coding
To ensure the reliability and consistency of the coding process, two expert coders initially met to collaboratively develop a comprehensive codebook consisting of ten codes: three parent codes and seven child codes. The full codebook can be seen in Appendix H.
After joint codebook creation, each coder independently coded all match errors produced by GPT 3.5TPE_25ep. Once the coding was complete, we computed inter-coder reliability and found a high level of agreement between the two coders with a Cohen’s kappa score of 0.87 (Cohen, 1960). The results of the qualitative coding are visualized in Figure 5 and detailed further below. Note that in the following text, parent codes from the codebook are italicized, while child codes are enclosed in quotation marks
@@ 5.4.2. Semantic Equivalence
The child code “Semantic Equivalence,” the sole child code of the parent code Semantic Equivalence, was the most prevalent code in our error analysis—accounting for 19/37 (51%) errors. The following two examples demonstrate this type of error. The text in quotation is the expert annotation, while the underlined sections are the model’s annotation:
@@
Figure 5. Breakdown by parent code of the various types of errors found from our qualitative analysis.
5.4.3. Incorrect Expert Annotations
The parent code Expert Labels Is Wrong had nine examples (24%). Occasional expert mis-annotations are expected for a task of this complexity. We are encouraged that there were relatively few examples under this parent code, supporting the quality of our ground truth. Importantly, for the examples in the “Expansive Ground Truth” (4/9, 44%) and “Partial Ground Truth” (3/9, 33%) child codes, the model performed the task more correctly than the expert annotator—either by omitting superfluous words included in the expert annotation or including necessary ones the expert annotator missed.
Consider the following example “Expansive Ground Truth” annotation. The text in quotation is the expert annotation, while the crossed out section is what the model correctly excluded:
@@ 5.4.3. Incorrect Expert Annotations
Conversely, the following example “Partial Ground Truth” annotations show the expert annotation in quotations, while the underlined words are what the model correctly choose to additionally include in its response:
@@
Combined, “Expansive Ground Truth” and “Partial Ground Truth” codes represent only 9 of the 37 errors (24%); however, we emphasize this finding as particularly exciting because they demonstrate that the model can identify precise annotations for the requested parameter. In other words, the model’s ability has surpassed that of our trained human annotators in these situations.
@@ 5.4.4. True Model Errors
The parent code Completion Errors accounted for 9 out of 37 coded examples (24%). Notably, a significant majority of these errors (6 out of 9, 67%) fall under the “Meaningful Subset” child code. A “Meaningful Subset“ annotation included a segment of the correct response, but missed words that altered the meaning of the annotation. In the following example, the text in quotation is the expert annotation while the underlined sections are the model’s annotation:
@@
Many of these completions do encompass enough of the correct answer for someone well-versed in GKC-CI to grasp the intended annotation. However, we consider these incomplete responses to be incorrect even though the model’s answer is a meaningful subset of the correct response.
Finally, the parent code Completion Is Wrong comprises the final 3 out of 9 codes (33%). All of these responses annotated text from within the sentence with no relation to the actual GKC-CI parameter. For example the model incorrectly annotates the underlined part of the following sentence as an attribute:
@@
Our qualitative analysis offers a comprehensive view of the model’s performance, detailing both its strengths and limitations. We find that appreciably more than half of the purported match errors are the codes “Semantic Equivalence,” “Expansive Ground Truth,” and “Partial Ground Truth.” These combined child codes constitute 27 examples, resulting in a $0.29$ percent increase in the number of correct annotations overall. This implies that our benchmarking metric of 90.65% model accuracy serves as a conservative estimate of model performance.
@@ 5.5. Training Data Review
Motivated by the nine expert annotation errors we uncovered in Section 5.4.3, we had two different CI experts review a random sample of 5% of our training data. These experts agreed with the training data on 90.57% of the sample, with the most common disagreement being the identification of occasional attribute parameters that were missed in the original annotations. The original annotators spent several months annotating thousands of parameters, so it is unsurprising that a few were missed. Furthermore, governing knowledge commons and contextual integrity are the topic of active research and subject to different interpretive nuances by different members of the community, which is reflected in different annotations. This training data review validates our results, as our best performing model’s 90.65% accuracy is close to the inter-expert annotation agreement, suggesting that the model approaches human performance on the task.
@@ 6. Limitations
Second, the LLMs we use do not have information about the input text outside of the context window. This is unlike a human lawyer, who could review the entire document while identifying information flow descriptions. Fortunately, the information needed to identify GKC-CI parameters typically appears in the text immediately surrounding the parameter itself. The decreased accuracy we observed when using a longer context window (Section 5.3.2, Appendix C) supports our use of one-sentence contexts.
Third, we do not perform coreference resolution, i.e., identifying when different sections of text refer to the same real world entity. For example, a sentence may refer to the website user multiple times, with each instance receiving a GKC-CI parameter annotation (e.g., “sender”). Although our model would accurately annotate each mention of the sender, it would not automatically recognize that these annotations refer to the same legal entity. If that were necessary for a downstream application, another model would be required. Since coreference resolution is an active research problem in its own right, and is independent from GKC-CI parameter annotation, we leave this for future development.
@@ 7. Example Applications
We applied our GPT 3.5TPE_25ep model to 456 privacy policies from the Princeton-Leuven Longitudinal Corpus of Privacy Policies (Amos et al., 2021) to demonstrate the type of analyses enabled by GKC-CI annotation at scale. This dataset contains over 1 million privacy policies from over 100,000 companies spanning more than two decades, making it an ideal data source. However, we note that the primary contribution of our project remains the LLM training and evaluation (Sections 45). This section is not meant to provide a comprehensive analysis of policies in the Princeton-Leuven dataset. Rather, we intend the following examples to inspire future work using our LLM annotation method and the annotated policies we provide. All 456 policies we annotated for the following analyses are publicly available at at the GitHub repository for this paper. [^8] https://github.com/JakeC007/Automated_GKC-CI_Privacy_Policy_Annotations
@@ 7.1. Longitudinal Privacy Policy Analysis
First, we chose 8 prominent companies and organizations [^9] Facebook, The New York Times, GitHub, Buzzfeed, Google, Bank of America, Electronic Frontier Foundation (EFF), and the National Science Foundation (NSF) representing a variety of sectors, including “big tech,” news, entertainment, finance, and government. This variety is useful because each sector has a unique approach to data collection, user engagement, and compliance with privacy regulations. Furthermore, these websites have undergone varying levels of public scrutiny. For instance, while Facebook and Google have faced major privacy debates, leading to numerous changes in their privacy policies, entities like nsf.gov operate under distinct governmental standards. This list also highlights geographic diversity concerning headquarters and user base, with some organizations primarily serving U.S. audiences while others have a global reach, necessitating compliance with various international privacy laws per the Brussels effect (Bradford, 2012).
For each of these companies and organizations, we used our model to annotate one privacy policy from every year that the company or organization appears in the dataset (128 policies total). The number of parameters in the policies of each company or organization over time are presented in Appendix I (Figure 17 and Tables 67).
@@ 7.2. Cross-Industry Privacy Policy Analysis
Figure 6 shows the fifteen policies with the highest variance of parameter type percentages. This includes a policy from apache.org with relatively few aim parameters and a policy from mozilla.org with relatively few attribute parameters. A quick review of the mozilla.org policy shows why – it defines all data covered by the policy in one sentence: “For us, ‘personal information’ means information which identifies you, like your name or email address.” The policy does not say which particular forms of personal information are connected to the descriptions of information use in the rest of the policy, leaving it ambiguous whether e.g., what specific customer data might be shared with other entities for “processing or providing products and services.” While this type of high-level analysis doesn’t necessarily imply the existence of policy ambiguities, it suggests that policies with high parameter type variance are promising candidates for a detailed evaluation through the lens of the GKC-CI framework. The parameter percentages for all 164 privacy policies are provided in Tables 89 in Appendix J.
@@
We next calculated the ratio of annotated GKC-CI parameters to the number of sentences in each policy. This provides a metric of the “density” of information transfer descriptions in the policy. Figure 7 shows these data for the 15 privacy policies with the highest ratio of annotated parameters to sentences. The ratios for all 164 privacy policies are provided in Tables 1011 in Appendix K. The top 15 policies include those from content distribution networks (b-cdn.net), web component frameworks (ampproject.org), Facebook, Microsoft (bing.com), Apple (icloud.com), Zoom, and social media websites (linkedin.com, snapchat.com, t.co, tumblr.com), among others.
@@ 7.3. GKC-CI Annotation Visualizer
We have developed a visualizer tool designed to enhance the interpretability of our LLM’s privacy policy annotations. This tool was created to empower researchers and others using our approach to better understand the complex data flows and governing principles that underpin privacy policies by visually distinguishing between different parameter tags. By highlighting various aspects of the policy in different colors, the visualizer makes the described information flows more accessible, allowing users to quickly grasp how their data may be used and protected under the terms of the policy.
The tool functions as a local Python script that processes both the raw text of the privacy policy and the output log from the LLM. It begins by matching specific text segments from the policy with those used in the LLM prompts. The visualizer then highlights these segments in different colors based on the associated annotations, making it easy for users to see which parts of the text correspond to specific GKC-CI parameters. This highlighted text is presented in a user-friendly GUI window, where users can scroll through and interact with the augmented policy. Additionally, it generates an output file that explains the highlighted segments, helping users understand the implications of the annotations. Note that if the LLM’s output has not match in the text we consider the output incorrect and do not include these instances in the visualization.
To illustrate the capabilities of the visualizer, we applied it to two versions of Facebook’s privacy policy from 2015 and 2019. Comparing visualized excerpts of these versions (Appendix L) effectively highlights how data practices and governing principles have evolved. For instance, there are more GKC-CI parameters in the 2019 excerpt than the 2015 excerpt, particularly more aims and consequences. The new policy does a better job informing users about the repercussions of the described data collection. The visualizer’s color-coded presentation makes it easier to understand how the policy has changed and will facilitate future research using GKC-CI annotation.
@@ 9. Conclusion
This paper demonstrates that high-accuracy annotation of contextual integrity (CI) and governing knowledge commons (GKC) parameters in privacy policies can be achieved using LLMs. We ultimately find that GPT 3.5TPE_25ep had the best performance, with an accuracy of 90.65% for exact string matches. While we find that proprietary LLMs outperformed open-source models, we report some valuable findings for researchers interested in performing LLM application studies. Namely, 1) that LLM size must be considered in context to model family; smaller, aligned models are the most economical choice, and 2) that library defaults are likely to introduce confounds and should be checked.
We demonstrate the usefulness of our fine-tuned model by annotating 456 privacy policies from the Princeton-Leuven Longitudinal Corpus of Privacy Policies (Amos et al., 2021). We show that large-scale GKC-CI annotation can be an effective tool for data exploration, highlighting changes in parameter frequency over time, policies with relatively high variances across parameter type percentages, and policies with relatively high ratios of parameters to sentences. This facilitates automated review of privacy policy updates to identify meaningful changes in information flow descriptions, potentially with normative implications. We have made our model training code, training and testing data, annotation visualizer, and all annotated policies publicly available [^10] https://github.com/JakeC007/Automated_GKC-CI_Privacy_Policy_Annotations in the hope that this work motivates future use of GKC-CI parameter annotation.
Acknowledgements.
@@ Appendix A Ground Truth Details
Table 3. Number of labeled parameters in ground-truth GKC-CI annotations of 16 privacy policies (rows) from popular websites and e-learning services. All of the labels from the manually-annotated privacy policies were combined together to create the ground-truth dataset used for evaluating all models. There is no difference in which labels were used for training/testing across different models.
@@ Appendix B Brat Annotation Legend
Figure 9. Legend in the customized brat annotation tool (Stenetorp et al., 2012) for expert annotators to use as reference.
Appendix C Performance On Sentence Versus Paragraph-Delineated Data As Epochs Vary
Figure 10. The solid blue line (“FT-Sentence Accuracy”) is the test-set accuracy of GPT 3.5TPE trained on sentence-delineated data at 1, 5, 10, 25, and 50 epochs. The dotted orange line (“FT-Paragraph Accuracy”) is the test-set accuracy of GPT 3.5TPE trained on paragraph-delineated data at 1, 5, 10, 25, and 50 epochs. An accurate annotation is one that is an exact string match.